Evaluate the following polynomial expression using Horner’s rule. Assume the value of x is 3.

𝑃(π‘₯)=10π‘₯τ°€ +6π‘₯τ° +3π‘₯τ°‚ βˆ’6π‘₯τ°ƒ +8π‘₯+15

Show all the intermediate steps.


1
Expert's answer
2021-05-07T14:23:26-0400


Problem is 𝑃(π‘₯) = 10x5Β + 6x4+ 3x3Β βˆ’ 6x2Β + 8π‘₯ + 15

At x= 3

Or "x_o=3"


So

This is of this type


f(x) = a0Β + a1x + a2x2Β + a3x3Β + a4x4Β + a5x5


Can be arranged as follows

at "x_o"

f(x0) = a0Β + x0(a1Β + x0(a2Β + x0(a3Β + x0(a4Β + a5x0))))


So

At k= 5

b5= a5=10


At k=4

b4=a4+xob5= 6+3Γ—10=36


At k=3

b3=a3+xob4=3+3Γ—36=111


At k=2

b2=a2+xob3=-6+3Γ—111=327


At k=1

b1=a1+xob2=8+3Γ—327=989


At k=0

b0=a0+xob1=15+3Γ—989=2982


Therefore

"\boxed{f(3)=2982}answer"